# VGC: A High-Performance Zone-Based Garbage Collector Architecture for Python with Partitioning and Parallel Execution

### v1

@@ 9 Related Work
Memory management and parallel execution have long been central challenges in high-level language runtimes. Python employs reference counting complemented by a cyclic garbage collector, prioritizing deterministic memory reclamation over parallel execution. While this design simplifies object lifetime tracking, it inherently limits CPU-bound workloads due to the Global Interpreter Lock (GIL).
Several efforts aim to mitigate these limitations, including Possible spelling mistake found., per-interpreter GIL proposals, and multiprocessing-based parallelism. However, these approaches introduce additional complexity in memory sharing, inter-process communication, or object ownership semantics.
Alternative runtimes such as PyPI and Gradually explore tracing garbage collection and just-in-time compilation to improve execution throughput. These systems focus primarily on speculative optimization and execution speed, often trading determinism, predictability, or memory stability for performance gains.
Parallel execution frameworks and task schedulers enable concurrency at the application level but remain external to runtime memory management, leaving object lifecycle control and garbage collection behavior largely unchanged.
In contrast, this work proposes a runtime architecture that explicitly separates execution partitioning (PPE), memory lifecycle control (VGC), and ephemeral execution handling (Yield Memory). Rather than relying on generational promotion, object migration, or speculative optimization, the proposed approach emphasizes deterministic behavior, bound memory usage, and explicit architectural control over parallel execution and object lifecycles.
@@ 10 Discussion
The experimental results validate the architectural principles underlying Partition and Parallel Execution (PPE), Virtual Garbage Collector VGC, and Yield Memory[5], demonstrate that CPU-bound workloads such as loops, recursion, deep recursion, and matrix operations can scale effectively across available cores without global synchronization or execution serialization.
@@
The VGC experiments confirm that the use of memory remains strictly bound under sustained allocation pressure. Objects do not migrate between zones; instead, they expire and are recreated on the basis of checkpoint evaluation, reinforcing the logical lifecycle model proposed in this work. This behavior contrasts with traditional generational garbage collectors that rely on object promotion and relocation.
@@
The checkpoint system enables a deterministic liveness evaluation in constant-time using compact bit fields, allowing concurrent execution without global locks or stop-the-world pauses. Yield Memory further complements this design by providing a lightweight execution path for tiny, short-lived operations, bypassing unnecessary memory management overhead while remaining compatible with the broader lifecycle system when persistence is required.
@@
Importantly, the observed performance and memory characteristics are achieved through architectural restructuring rather than aggressive optimization, speculative execution, or hardware acceleration, which highlights that scalability and predictability can be addressed at the runtime design level.
11 Future Goals
@@
A key direction for future work is the extension of VGC into a dual-layer model consisting of Active VGC and Passive VGC, each optimized for distinct phases of program execution.
@@
Active VGC is intended to operate at runtime, managing dynamic allocations, ephemeral objects, and adaptive lifecycle decisions under live execution conditions. This layer focuses on minimizing latency, reducing pause behavior, and responding to real-time access patterns.
@@
Passive VGC, in contrast, is envisioned as a compile-time or pre-execution analysis layer. It would handle static allocations, predictable object lifetimes, and structural memory planning based on program analysis, enabling further reduction of runtime overhead.
@@
In this model, Active and Passive VGC layers are logically separated but coordinated through shared metadata and checkpoint semantics. The present work does not implement or evaluate this dual-layer design; it is proposed as a future architectural evolution of VGC aimed at improving scalability, predictability, and integration with ahead-of-time optimization pipelines.
This work focuses on validating architectural feasibility rather than delivering a production-ready runtime Several directions remain open for future exploration.
@@
Integration with existing language runtimes, particularly Python, represents a natural next step It includes evaluating compatibility with Python’s object model, the C-extension ecosystem and interpreter internals.
@@
Extending PPE beyond dual-core execution to higher core counts and heterogeneous systems may provide further insight into scheduling strategies and partition granularity in various hardware configurations.
@@
Formal verification of checkpoint logic and zone lifecycle invariants could strengthen correctness guaranties, especially in highly concurrent environments.
@@
Finally, while this work intentionally excludes just-in-time compilation, ahead-of-time compilation, SIMD acceleration, and GPU offloading, future research may explore how such techniques could coexist with PPE, VGC, and Yield Memory without compromising determinism or bounded memory behavior.
@@ 12 Conclusion
This paper introduced a partition-aware runtime architecture (PPE), a virtual garbage collection model (VGC), and a Yield Memory execution layer designed to address fundamental limitations in parallel execution and memory predictability within high-level language runtimes.
By decoupling execution from memory lifecycle management, employing zone-local object expiration, and using checkpoint-based liveness evaluation, the proposed system achieves parallel scalability with strictly bounded memory usage. Yield Memory further optimizes ephemeral operations by avoiding unnecessary memory management overhead.
The experimental results demonstrate that predictable scalability and memory efficiency can be achieved through architectural design rather than speculative optimization or complex garbage collection heuristics. This work establishes a foundation for future exploration of deterministic, parallel, and memory-efficient runtime systems and is released as an open research platform to encourage further investigation and refinement by the community.

### v2

@@ 8.29 Checkpoint-Based Object Liveness and Zone Lifecycle Control
8.30 Benchmark Summary
The benchmark evaluation is divided into two independent components corresponding to Partitioned Parallel Execution (PPE) and the Virtual Garbage Collector (VGC). These components are evaluated separately to avoid conflating execution behavior with memory management behavior. The PPE benchmarks characterize execution behavior under CPU-bound workloads, including loop-based computation, recursive evaluation, deep recursion, and matrix multiplication. In single-core configurations without partitioning, execution time scales proportionally with workload size, establishing predictable baseline behavior. Under dual-core execution with two partitions, the results indicate reduced execution time relative to single-core baselines across the evaluated workloads. This behavior reflects reduced execution span achieved through static partitioning of execution ranges rather than speculative optimization or runtime scheduling heuristics. Variability across attempts remains limited, indicating stable execution behavior under the tested configurations.
Memory usage observed during PPE benchmarks remains bounded and exhibits predictable growth patterns aligned with partition size and recursion depth. In deep recursion workloads, increased memory consumption corresponds to intentional stack and allocation pressure introduced by recursion depth, rather than unbounded allocation growth or memory leakage.
The VGC benchmarks evaluate memory allocation behavior independently of execution partitioning. Under sustained allocation pressure, including balanced, imbalanced, and zone-skewed access patterns, the number of real allocations remains constant while object reuse dominates allocation requests. Objects expire and are reused within their assigned zones without cross-zone migration, demonstrating strict space boundedness and effective checkpoint-based lifecycle control.
These benchmarks are intended to evaluate architectural feasibility and behavioral characteristics rather than establish absolute performance superiority over production runtimes. The evaluation excludes just-in-time compilation, ahead-of-time optimization, SIMD acceleration, and GPU offloading. Consequently, the results should be interpreted as evidence that partition-aware execution and lifecycle-driven memory management can independently provide predictable execution behavior and bounded memory usage at the architectural level.
@@ 9 Related Work
Memory management and parallel execution have long been central challenges in high-level language runtimes. The Python interpreter employs reference counting complemented by a cyclic garbage collector, prioritizing deterministic memory reclamation and predictable object lifetime semantics. While this design simplifies memory management, it restricts scalable parallel execution for CPU-bound workloads due to the presence of the Global Interpreter Lock (GIL).
Several approaches have been proposed to mitigate these limitations, including multiprocessing-based parallelism, per-interpreter GIL designs, and subinterpreter isolation mechanisms. Although these techniques improve concurrency, they introduce additional complexity related to memory sharing, inter-process communication, and object ownership semantics, often shifting the burden of coordination to the application layer.
Alternative Python runtimes, such as PyPy and GraalPy, explore tracing garbage collection and just-in-time compilation to improve execution throughput. These systems emphasize speculative optimization and dynamic code generation, frequently trading deterministic execution behavior, memory predictability, or stable object lifetimes for performance gains.
Parallel execution frameworks and task schedulers further enable concurrency at the application level; however, they remain external to runtime memory management. As a result, object lifecycle control, garbage collection behavior, and allocation strategies remain largely unchanged, limiting their ability to address memory-related bottlenecks holistically.
In contrast, this work proposes a runtime-oriented architecture that explicitly separates execution partitioning through Partitioned Parallel Execution (PPE), memory lifecycle control via the Virtual Garbage Collector (VGC), and ephemeral execution handling using Yield Memory. Rather than relying on generational promotion, object migration, or speculative optimization, the proposed approach emphasizes deterministic behavior, bounded memory usage, and explicit architectural control over parallel execution and object lifecycles.
@@ 10 Discussion
The experimental results support the architectural principles underlying Partitioned Parallel Execution (PPE), the Virtual Garbage Collector (VGC), and Yield Memory [5]. The evaluated workloads indicate that CPU-bound execution patterns, including loop-based computation, recursive functions, and matrix-oriented operations, can be structured to exploit available cores without introducing global synchronization or execution serialization at the architectural level.
@@
The VGC experiments demonstrate that memory usage remains bounded under sustained allocation pressure. Objects do not migrate between zones; instead, they expire and are recreated based on checkpoint evaluation, consistent with the logical lifecycle model proposed in this work. This behavior differs from traditional generational garbage collectors, which rely on object promotion and relocation across memory regions.
@@
The checkpoint system enables deterministic liveness evaluation using compact bitfields, allowing object state assessment with constant-time complexity. This design supports concurrent execution without requiring global locks or stop-the-world pauses. Yield Memory further complements the architecture by providing a lightweight execution path for short-lived operations, reducing interaction with memory management mechanisms while remaining compatible with the broader lifecycle model when persistence is required.
@@
Overall, the observed execution and memory characteristics arise from architectural decomposition and explicit lifecycle control rather than aggressive optimization techniques, speculative execution, or hardware-specific acceleration. These results suggest that improvements in scalability and predictability can be achieved through runtime-level design choices.
11 Future Work
@@
A key direction for future research is the extension of the Virtual Garbage Collector (VGC) into a dual-layer architecture consisting of an Active VGC and a Passive VGC, each optimized for distinct phases of program execution.
@@
The Active VGC is intended to operate at runtime, managing dynamic allocations, ephemeral objects, and adaptive lifecycle decisions under live execution conditions. This layer focuses on minimizing latency, reducing pause behavior, and responding to runtime access patterns.
@@
In contrast, the Passive VGC is envisioned as a compile-time or pre-execution analysis layer. It would manage static allocations, predictable object lifetimes, and structural memory planning derived from program analysis, with the goal of reducing runtime memory management overhead.
@@
In this model, the Active and Passive VGC layers are logically separated but coordinated through shared metadata and checkpoint semantics. The present work does not implement or evaluate this dual-layer design; it is proposed as a potential architectural evolution of VGC aimed at improving scalability, predictability, and integration with ahead-of-time optimization pipelines.
@@
The current study focuses on validating architectural feasibility rather than delivering a production-ready runtime. Several additional directions remain open for future exploration. Integration with existing language runtimes, particularly Python, represents a natural next step and includes evaluating compatibility with Python’s object model, interpreter internals, and the C-extension ecosystem.
@@
Extending Partitioned Parallel Execution (PPE) beyond dual-core execution to higher core counts and heterogeneous systems may provide further insight into scheduling strategies and partition granularity across diverse hardware configurations.
@@
Formal verification of checkpoint logic and zone lifecycle invariants could further strengthen correctness guarantees, particularly in highly concurrent execution environments.
@@
Finally, although this work intentionally excludes just-in-time compilation, ahead-of-time compilation, SIMD acceleration, and GPU offloading, future research may explore how such techniques could coexist with PPE, VGC, and Yield Memory without compromising determinism or bounded memory behavior.
@@ 12 Conclusion
This paper presented a partition-aware runtime architecture based on Partitioned Parallel Execution (PPE), a Virtual Garbage Collector (VGC), and a Yield Memory execution layer, with the goal of addressing challenges related to parallel execution structure and memory predictability in high-level language runtimes. By decoupling execution partitioning from memory lifecycle management, employing zone-local object expiration, and utilizing checkpoint-based liveness evaluation, the proposed architecture provides a structured approach to managing execution and memory behavior. Yield Memory complements this design by supporting lightweight handling of short-lived operations, reducing unnecessary interaction with memory management mechanisms. The experimental evaluation indicates that improved scalability characteristics and bounded memory behavior can be achieved through architectural decomposition and explicit lifecycle control, rather than reliance on speculative optimization or complex garbage collection heuristics. While the system remains an experimental prototype, the results suggest that runtime-level design choices can play a significant role in improving predictability and scalability. This work establishes a foundation for further investigation into deterministic and partition-aware runtime architectures. The implementation and design are released as an open research platform to facilitate continued experimentation, validation, and refinement by the research community.